Rachel just received her transcript showing her grades in five academic classes. In how many ways could she have earned "A's" in at least three of her classes?
Answer by stanbon(75887)

Rachel just received her transcript showing her grades in five academic classes. In how many ways could she have earned "A's" in at least three of her classes?
------
Ans:
5C3 + 5C4+ 5C5
5C3 = (5*4)/(1*2) = 10 ways
(5C4) = 5
(5C5) = 1
Total = 10+5+1 = 16
====================
Cheers,
Stan H.
===========